Consider the combustion products of decane (C_10H_
Answer #1

a)
Equivalence ratio = 1.25 = Mass of 1 mole Air / Mass of 1 mole Decane
So, Mass of 1 mole air = 1.25 * 142 = 177.5 gm/mol
Number of moles of air = 177.5 / 29 = 6.12 moles
Number of moles of Oxysgen = 21% = 21% of 6.12 = 1.224 moles

Reaction is :
C10H22 + 31/2 O2 --> 10CO2 + 11H2O
1 - x 1.224 - 31/2 * x 10*x 11*x
If complete combustion happens then some Decane will be left,
x = 1.224 / 31 = 0.079 moles
Decane left = 1 - 0.079 =  0.921 moles
CO2 formed = 0.79 moles
H2O formed = 0.869 moles